OpenStudy (anonymous):

Easy question: Sn = 1[2(3)+(4-1)

OpenStudy (anonymous):

\[\Large \text{\S}_n=\frac{1[2\cdot 3+(4-1)\cdot 5]}{2} \] ?? you mean like this ?

OpenStudy (anonymous):

\[\Large \text{\S}_n \frac{6+15}{2}=\frac{21}{2}=10.5\]

OpenStudy (anonymous):

is this sum of arithmetic sequences ??